OpenStudy (diamond77108):

WILL MEDAL 1.4t−0.4(t−3.1)=5.8

OpenStudy (studygurl14):

Solve for t? First distribute the -0.4

OpenStudy (diamond77108):

1.4t-0.4t- 1.24=5.8

OpenStudy (diamond77108):

t-1.24=5.8

OpenStudy (studygurl14):

Almost. t + 1.24 = 5.8 keep going

OpenStudy (diamond77108):

t=4.56

OpenStudy (studygurl14):

yes